# Computing Evolutionarily Stable Strategies in Imperfect-Information Games

### v1

@@ 1 Introduction
There exists a polynomial-time algorithm for computing Nash equilibrium (NE) in two-player zero-sum normal-form games, while for two-player non-zero-sum and multiplayer games computing an NE is PPAD-complete and it is widely conjectured that no efficient (polynomial-time) algorithm exists. However, several algorithms have been devised that perform well in practice. The problem of computing whether a game has an ESS was shown to be both NP-hard and CO-NP hard and also to be contained in $\Sigma^{P}_{2}$ (the class of decision problems that can be solved in nondeterministic polynomial time given access to an NP oracle) Etessami and Lochbihler (2008). Subsequently it was shown that the exact complexity of this problem is that it is $\Sigma^{P}_{2}$-complete Conitzer (2013), and even more challenging for more than two players Blanc and Hansen (2021). Note that this result is for determining whether an ESS exists (as discussed above there exist games which have no ESS), not for the complexity of computing an ESS in games for which one exists. Thus, computing an ESS is significantly more difficult than computing an NE, which is not surprising since it is a refinement. Several approaches have been proposed for computing ESS in two-player normal-form games Haigh (1975); Abakuks (1980); Broom and Rychtář (2013); Bomze (1992); McNamara et al. (1997), and recently an algorithm has been developed for multiplayer normal-form games Ganzfried (2025a). However, these algorithms are based no support enumeration and are not applicable to extensive-form games.
@@ 2 Evolutionarily Stable Strategies
The final two constraints are called complementarity slackness conditions (CSC), and the full system is known as a linear complementarity problem (LCP). This LCP can be solved using Lemke’s algorithm Lemke (1965) or the related Lemke-Howson algorithm Lemke and Howson (1964).
@@
for every terminal history $z$, where $\pi$ is the player-swapping automorphism of the game tree. Under symmetry, we may identify a behavioral strategy with its realization plan $\mathbf{x}$ in sequence form, subject to the information-set consistency constraints $\mathbf{E}\mathbf{x}=\mathbf{e}$ and $\mathbf{x}\geq\mathbf{0}$. Let $\boldsymbol{\sigma}$ denote a behavioral strategy, and let $u(\boldsymbol{\sigma},\boldsymbol{\tau})$ be the expected payoff to a player who plays $\boldsymbol{\sigma}$ against an opponent playing $\boldsymbol{\tau}$. Because the game is symmetric, the classical definition of evolutionary stability applies directly to behavioral strategies.
@@ 4 Experiments
Our algorithm identifies seven symmetric Nash equilibria. In every equilibrium, players assign probability 1 to the proliferative phenotype P at the “fav” information set, which reflects that proliferation is optimal when the signal indicates likely low therapy. The equilibria differ only in their behavior at the “unfav” information set. Here the QCP in Algorithm 3 uncovers a mixture of pure and mixed SNE that distribute probability across $\{P,R,Q\}$ following an unfavorable signal. Several of these arise from degeneracy: under high therapy, the resistant R and quiescent Q phenotypes yield nearly identical continuation values, producing a continuum of payoff-equivalent best responses. Table 3 summarizes all seven SNE, and the full computation required only 0.287 seconds.
@@ 5 Conclusion
Our algorithm is fundamentally different from approaches for computing ESS in normal-form games. Those approaches are based on enumerating supports, and for each support computing a potential SNE and ESS. While the number of supports is exponential in the number of pure strategies, for relatively small games it can be done efficiently in practice. By contrast, it does not make sense to think of the concept of supports in extensive-form games, as enumeration of all of them is infeasible. We also note that several efficient preprocessing procedures developed for the normal-form setting are no longer applicable to the extensive form. A recent approach applies a “strict NE shortcut” procedure that immediately classifies an SNE as an ESS if it satisfies a simple property ensuring that it is a strict equilibrium Ganzfried (2025a). That approach involves testing whether there is a unique best response to the SNE, which is simple to test in normal-form games but not in extensive-form games. That work also applies a procedure that discards SNE that fail a “pure mutant screen”, which is also easy to do in normal-form games but not in extensive-form games. By contrast our approach creates a QCPC that searches for potential pure and mixed mutants. Perhaps improved preprocessing procedures can be developed that are applicable to extensive-form games and can reduce the number of QCPC solves needed by Algorithm 4.

### v2

@@ 1 Introduction
There exists a polynomial-time algorithm for computing Nash equilibrium (NE) in two-player zero-sum normal-form games, while for two-player non-zero-sum and multiplayer games computing an NE is PPAD-complete and it is widely conjectured that no efficient (polynomial-time) algorithm exists. However, several algorithms have been devised that perform well in practice. The problem of computing whether a game has an ESS was shown to be both NP-hard and CO-NP hard and also to be contained in $\Sigma^{P}_{2}$ (the class of decision problems that can be solved in nondeterministic polynomial time given access to an NP oracle) Etessami and Lochbihler (2008). Subsequently it was shown that the exact complexity of this problem is that it is $\Sigma^{P}_{2}$-complete Conitzer (2013), and even more challenging for more than two players Blanc and Hansen (2021). Note that this result is for determining whether an ESS exists (as discussed above there exist games which have no ESS), not for the complexity of computing an ESS in games for which one exists. Thus, computing an ESS is significantly more difficult than computing an NE, which is not surprising since it is a refinement. Several approaches have been proposed for computing ESS in two-player normal-form games Haigh (1975); Abakuks (1980); Broom and Rychtář (2013); Bomze (1992); McNamara et al. (1997), and recently an algorithm has been developed for multiplayer normal-form games Ganzfried (2025a). However, these algorithms are based on support enumeration and are not applicable to extensive-form games.
@@ 2 Evolutionarily Stable Strategies
The final two constraints are called complementarity slackness conditions, and the full system is known as a linear complementarity problem (LCP). This LCP can be solved using Lemke’s algorithm Lemke (1965) or the related Lemke-Howson algorithm Lemke and Howson (1964).
@@
for every terminal history $z$, where $\pi$ is the player-swapping automorphism of the game tree. Under symmetry, we may identify a behavioral strategy with its realization plan $\mathbf{x}$ in sequence form, subject to the information-set consistency constraints $\mathbf{E}\mathbf{x}=\mathbf{e}$, $\mathbf{x}\geq\mathbf{0}.$ Let $\boldsymbol{\sigma}$ denote a behavioral strategy, and let $u(\boldsymbol{\sigma},\boldsymbol{\tau})$ be the expected payoff of playing $\boldsymbol{\sigma}$ against an opponent playing $\boldsymbol{\tau}$. Because the game is symmetric, the classical definition of evolutionary stability applies directly to behavioral strategies, with $\mathbf{B}=\mathbf{A}^{\top},$ $\mathbf{F}=\mathbf{E},$ $\mathbf{f}=\mathbf{e}.$
@@ 4 Experiments
Our algorithm identifies seven symmetric Nash equilibria. In every equilibrium, players assign probability 1 to the proliferative phenotype P at the “fav” information set, which reflects that proliferation is optimal when the signal indicates likely low therapy. The equilibria differ only in their behavior at the “unfav” information set. Here the QCP in Algorithm 3 uncovers a mixture of pure and mixed SNE that distribute probability across $\{P,R,Q\}$ following an unfavorable signal. Several of these arise from degeneracy: under high therapy, the resistant R and quiescent Q phenotypes yield nearly identical continuation values, producing a continuum of payoff-equivalent best responses. Table 3 summarizes all seven SNE. The full computation required only 0.287 seconds.
@@ 5 Conclusion
Our algorithm is fundamentally different from approaches for computing ESS in normal-form games. Those approaches are based on enumerating supports, and for each support computing a potential SNE and ESS. While the number of supports is exponential in the number of pure strategies, for relatively small games it can be done efficiently in practice. By contrast, it does not make sense to think of the concept of supports in extensive-form games, as enumeration of all of them is infeasible. We also note that several efficient preprocessing procedures developed for the normal-form setting are no longer applicable to the extensive form. A recent approach applies a “strict NE shortcut” procedure that immediately classifies an SNE as an ESS if it satisfies a simple property ensuring that it is a strict equilibrium Ganzfried (2025a). That approach involves testing whether there is a unique best response to the SNE, which is simple to test in normal-form games but not in extensive-form games. That work also applies a procedure that discards SNE that fail a “pure mutant screen,” which is also easy to do in normal-form games but not in extensive-form games. By contrast our approach creates a QCPC that searches for potential pure and mixed mutants. Perhaps improved preprocessing procedures can be developed that are applicable to extensive-form games and can reduce the number of QCPC solves needed by Algorithm 4.